# Melting of quantum Hall Wigner and bubble crystals

### v1

@@ Abstract
A two-dimensional crystal melts via the proliferation and unbinding of topological defects, yet quantitatively predicting the melting temperature $T_{m}$ in real systems is challenging. Here we resolve this discrepancy in quantum Hall electron bubble phases by combining Corbino-geometry transport experiment in an ultraclean GaAs/AlGaAs quantum well for Landau levels $2\sim 5$ with Hartree–Fock elasticity and the full Kosterlitz–Thouless–Halperin–Nelson–Young melting criterion including the finite-temperature renormalization-group calculation. The theoretically obtained $T_{m}$ quantitatively captures the measured solid-liquid phase transition boundaries across all probed ranges, validating the bubble-crystal interpretation and establishing defect-mediated melting as a predictive framework for strongly interacting electronic solids. This agreement further supports using bulk transport to probe the energetics of topological defects and screening in quantum Hall physics, and the approach is readily extendable to other electronic crystals, including the generalized Wigner crystal in moiré Chern bands.
@@
Introduction– Although melting is a ubiquitous phenomenon, the melting points of only a small portion of solids have been accurately measured due to rigid experimental requirements, and reliable theoretical predictions are often nontrivial. The challenge becomes particularly acute at ultralow temperatures, where melting is increasingly dominated by quantum fluctuations rather than thermal phonons. In two dimensions (2D), the situation is further constrained fundamentally; there, true long-range crystalline order does not exist at any finite temperature due to the Mermin-Wagner theorem [34]. Kosterlitz-Thouless-Halperin-Nelson-Young (KTHNY) theory establishes that 2D melting is a continuous phase transition driven by the proliferation and unbinding of topological defects [21, 19, 38, 47, 12]. Its quantitative predictions have been verified in classical Wigner crystals (WC), where the melting temperature $T_{m}$ exhibits the universal KTHNY scaling $T_{m}\propto\sqrt{n}$ ($n$ is the electron’s density) with electron-electron interaction being Coulomb [18, 43, 13, 37]. However, Wigner solids in strong magnetic fields at ultralow temperatures violate this scaling and display a nontrivial density dependence of $T_{m}$ [4]. In these strongly correlated systems, the lattice is largely softened by the strong quantum and thermal fluctuations, resulting in the overestimation of stiffness and $T_{m}$. Recent technical advances in van der Waals materials have facilitated the direct imaging of 2D electron solids and revealed the intertwined fluctuations and phase competition in the system, indicated by the short-range crystallinity and pronounced lattice softening [32, 30, 29, 44]. These observations underscore the complexity of quantum electronic lattices.
@@
The 2D electron bubble phases in ultra-high-mobility GaAs heterostructures provide an ideal platform for investigating these issues. Like the conventional Wigner crystal, a bubble phase is an interaction-driven electron solid, but features a generalized lattice where multiple electrons could cluster within one unit cell, which is predicted to be the ground state of a two dimensional electron gas (2DEG) under a perpendicular magnetic field at high Landau levels (LLs) [16, 24, 36, 41, 46, 27, 26, 9, 45, 15, 5, 3, 42]. Recent scanning probe experiments have directly imaged the analogous multi-electron crystalline phases in twisted WS2 moiré superlattices and manifested their rich real-space density profiles [31]. Previous theoretical studies of Wigner and bubble crystals have mainly focused on zero-temperature energetics based on Hartree-Fock (HF) or density-matrix renormalization group (DMRG) calculations [33, 25, 6, 28, 8, 11, 7], which capture the main features of phase competition and phase diagrams, but do not provide quantitative predictions for $T_{m}$. Predicting melting in 2D electron bubble phases is not merely a matter of computational complexity, but requires understanding and careful treatment of collective excitations coupled with Landau-level quantization and many-body interactions. Particularly in high LLs, the electron wavefunctions are strongly reshaped by Landau-level form factors, and the resulting spatial separation significantly weakens the short-range Coulomb repulsion and stabilizes charge clustering. Additionally, the presence of impurity potentials in real materials breaks the quasi-long-range order of 2D crystals, meaning the reliability of theoretical calculations must be evaluated through direct comparison with experimental data.
@@
In this Letter, we combine systematic transport measurements of electron bubble phases in high-mobility GaAs/AlGaAs 2DEGs (from LLs $N=2$ to $5$) with theoretical calculations to quantitatively capture the melting process. By applying LL-projected Hartree–Fock elasticity with the full KTHNY melting criterion and its renormalization-group (RG) flow, we for the first time obtained melting temperatures $T_{m}$ that quantitatively agree with experiment for the bubble phases across different LLs and spin branches. Our results established defect-mediated topological melting as a predictive framework for strongly correlated quantum Hall electron solids, accurately capturing the underlying connection between microscopic LL physics and macroscopic finite-temperature phase boundaries. In addition, we propose that the experimental measurement of the solid-liquid transition provides a quantitatively reliable bulk probe of screening and defect physics in quantum Hall solids in the presence of strong electron-electron interaction.
@@
Fig. 1a shows the longitudinal conductivity $\sigma$ of the Corbino-geometry sample (in units of the quantum conductance, $e^{2}/h$) measured at $40\,\text{mK}$ as a function of magnetic field $B$ with corresponding filling factors ($\nu=nh/eB$) indicated on the top axis. The data is dominated by large conductance peaks centered near half-integer filling factors ($i+1/2$), which originate from emerging stripe-like charge order. The electron bubble phases of interest are identified by the conductivity minima located between these major features and adjacent satellite conductance peaks. For clarity, we adopt the conventional notation [10] $Na$ and $Nd$ (e.g., $4a,4d$) to label different bubble phases, where the integer $N$ indexes the specific spin-split Landau level branch, and the suffixes $a/d$ denote particle-hole conjugate states within that branch. To probe the evolution and thermal stability of these states, we track the change of magneto-conductivity at different temperatures. We find that the broad conductivity minima initially observed at base temperature gradually narrows and vanishes at a certain point. This evolution marks the melting of electron-bubble phases, allowing the extraction of the solid-liquid transition (SLT) temperature ($T_{m}$) by identifying peaks in the non-monotonic temperature dependence of the conductivity. We also observe that the thermal stability of bubble phases varies significantly with filling factor; $T_{m}$ is maximized at an optimal filling factor (indicated by the red triangle in Fig. 1b) and decreases away from this point, forming a characteristic dome-like phase diagram.
@@
As a large melting temperature indicates a more stable phase, we fixed the magnetic field at these filling factors and performed a detailed temperature-dependent conductivity measurement (Fig. 1c). The melting temperature ($T_{m}$) from conductivity peaks is extracted as a function of the partial filling factor ($\nu=\nu_{0}-2N$) at different Landau levels (Fig. 2, top panel). With increasing $\nu$ in a given Landau level, $T_{m}$ roughly follows a smooth decreasing trend, yet exhibits a discrete jump when transitioning to a higher Landau level. This behavior can be qualitatively explained by adapting the universal KTHNY transition picture to the bubble phase. If we treat the bubbles as classical objects with effective charge $Q=Me$ (where $M$ is the number of electrons per bubble), the KTHNY theory yields the melting criterion [17] $4\pi k_{B}T_{c}=0.283e^{2.5}M^{1.5}\sqrt{\frac{B\nu^{*}}{h}}$. Where $\nu^{*}=\nu\text{ mod 1/2}$ is the partial filling factor defined as the decimal part of filling factors. This equation shows the transition temperature is a function of site electron occupation M, which is theoretically predicted to be dependent on Landau level $N$, with $M=N$ and $M=N+1$ being the most possible configurations. Assuming M=N under the classical picture, we calculate the expected temperature ratios between consecutive Landau levels and compare them with our experimental results. The predicted ratios of 1.49 (for $N=2\to 3$) and 1.07 (for $N=3\to 4$) show reasonable agreement with the observed values of 1.35 and 1.17, respectively.
@@
However, despite capturing the qualitative trend of the inter-level jumps, this simple classical picture remains fundamentally inadequate. The absolute melting temperatures derived from the formula are one order of magnitude lower than experimental values. Moreover, the functional dependence within a Landau level deviates significantly from the prediction. Our data exhibits a nearly linear dependence of $T_{m}$ on the magnetic field, while classical theory predicts a square-root scaling ($\propto\sqrt{B}$).
@@
In conclusion, we perform bulk Corbino transport measurements that track the evolution of electron bubble phases across a wide range of Landau levels, and develop a theoretical framework that quantitatively captures their melting temperatures. Beyond identifying the conductivity and thus $T_{m}$ of a 2D electronic crystal, the comparison between experiment and theory allows us to extract the dislocation-core parameter $\alpha$ and the screening strength $\gamma$ in the relevant Landau level. Here $\alpha$ quantifies the relative importance of dislocation core energetics versus long-wavelength elasticity (i.e. the fragility of the electronic solid to fluctuation-driven phase transition), while $\gamma$ calibrates the effective interaction, thereby constraining the competing electronic phases as well as the energy scales of their collective modes and excitations. The experimentally obtained SLT boundary therefore provides a bulk constraint on the microscopic cost of creating topological defects and the dressing of interaction.
Our strategy can be extended to include disorder, anisotropy, finite thickness, and Landau-level mixing. More broadly, it provides a route to study the melting and phase competition in other 2D electronic crystals, including solids proximate to fractional quantum Hall liquids, where the fundmental effective particles are emergent quasiparticles (anyons). Understanding the existence of solid phases there is essential for studying the boundaries of topological order. It also applies to the recently observed generalized Wigner and multi-electron crystals in TMD moiré systems [30, 39, 20, 31], where the required energetics and elastic response can be computed from band-projected form factors and interactions, enabling a unified description of flat-band screening, interaction dressing, and the stability of charge-ordered phases in Chern-band systems.

### v2

@@ Abstract
A two-dimensional crystal melts via the proliferation and unbinding of topological defects, yet quantitatively predicting the melting temperature $T_{m}$ in real systems is challenging. Here we resolve this discrepancy in quantum Hall electron bubble phases by combining Corbino-geometry transport experiment in an ultraclean GaAs/AlGaAs quantum well for Landau levels $2\sim 5$ with Hartree–Fock elasticity and the full Kosterlitz–Thouless–Halperin–Nelson–Young (KTHNY) melting criterion including the finite-temperature renormalization-group calculation. The theoretically obtained $T_{m}$ quantitatively captures the measured solid-liquid phase transition boundaries across all probed ranges, validating the bubble-crystal interpretation and establishing defect-mediated melting as a predictive framework for strongly interacting electronic solids. This agreement further supports using bulk transport to probe the energetics of topological defects and screening in quantum Hall physics, and the approach is readily extendable to other electronic crystals, including the generalized Wigner crystal in moiré Chern bands.
@@
Introduction– Although melting is a ubiquitous phenomenon, the melting points of only a small portion of solids have been accurately measured due to rigid experimental requirements, and reliable theoretical predictions are often nontrivial. The challenge becomes particularly acute at ultralow temperatures, where melting is increasingly dominated by quantum fluctuations rather than thermal phonons. In two dimensions (2D), the situation is further constrained fundamentally; there, true long-range crystalline order does not exist at any finite temperature due to the Mermin-Wagner theorem [34]. Kosterlitz-Thouless-Halperin-Nelson-Young (KTHNY) theory establishes that 2D melting is a continuous phase transition driven by the proliferation and unbinding of topological defects [22, 20, 38, 51, 13]. Its quantitative predictions have been verified in classical Wigner crystals (WC), where the melting temperature $T_{m}$ exhibits the universal KTHNY scaling $T_{m}\propto\sqrt{n}$ ($n$ is the electron’s density) with electron-electron interaction being Coulomb [19, 47, 14, 37]. However, Wigner solids in strong magnetic fields at ultralow temperatures violate this scaling and display a nontrivial density dependence of $T_{m}$ [4, 10, 11, 45]. In these strongly correlated systems, the lattice is largely softened by the strong quantum and thermal fluctuations, resulting in the overestimation of stiffness and $T_{m}$. Recent technical advances in van der Waals materials have facilitated the direct imaging of 2D electron solids and revealed the intertwined fluctuations and phase competition in the system, indicated by the short-range crystallinity and pronounced lattice softening [32, 30, 29, 48]. These observations underscore the complexity of quantum electronic lattices.
@@
The 2D electron bubble phases in ultra-high-mobility GaAs heterostructures provide an ideal platform for investigating these issues. Like the conventional Wigner crystal, a bubble phase is an interaction-driven electron solid, but features a generalized lattice where multiple electrons could cluster within one unit cell, which is predicted to be the ground state of a two dimensional electron gas (2DEG) under a perpendicular magnetic field at high Landau levels (LLs) [36, 44, 26, 10, 11, 43, 40, 42, 27, 9, 49, 16, 3, 46]. Recent scanning probe experiments have directly imaged the analogous multi-electron crystalline phases in twisted WS2 moiré superlattices and manifested their rich real-space density profiles [31]. Previous theoretical studies of Wigner and bubble crystals have mainly focused on zero-temperature energetics based on Hartree-Fock (HF) or density-matrix renormalization group (DMRG) calculations [33, 25, 6, 5, 50, 28, 8, 12, 7], which capture the main features of phase competition and phase diagrams, but do not provide quantitative predictions for $T_{m}$. Predicting melting in 2D electron bubble phases is not merely a matter of computational complexity, but requires understanding and careful treatment of collective excitations coupled with Landau-level quantization and many-body interactions. Particularly in high LLs, the electron wavefunctions are strongly reshaped by Landau-level form factors, and the resulting spatial separation significantly weakens the short-range Coulomb repulsion and stabilizes charge clustering. Additionally, the presence of impurity potentials in real materials breaks the quasi-long-range order of 2D crystals, meaning the reliability of theoretical calculations must be evaluated through direct comparison with experimental data.
@@
In this Letter, we combine systematic transport measurements of electron bubble phases in high-mobility GaAs/AlGaAs 2DEGs (from LLs $N=2$ to $5$) with theoretical calculations to quantitatively capture the melting process. By applying LL-projected Hartree-Fock elasticity with the full KTHNY melting criterion and its renormalization-group (RG) flow, we for the first time obtained melting temperatures $T_{m}$ that quantitatively agree with experiment for the bubble phases across different LLs and spin branches. Our results established defect-mediated topological melting as a predictive framework for correlated quantum Hall electron solids, accurately capturing the underlying connection between microscopic LL physics and macroscopic finite-temperature phase boundaries. In addition, comparison with experiments enables the extraction of topological defect energetics and interaction screening, making the measurement of the solid-liquid transition (SLT) a quantitatively reliable bulk probe of screening and defect physics in quantum Hall solids in the presence of strong electron-electron interaction.
@@
Fig. 1a shows the longitudinal conductivity $\sigma$ of the Corbino-geometry sample (in units of the quantum conductance, $e^{2}/h$) measured at $40\,\text{mK}$ as a function of magnetic field $B$ with corresponding filling factors ($\nu=nh/eB$) indicated on the top axis. The data is dominated by large conductance peaks centered near half-integer filling factors ($i+1/2$), which originate from emerging stripe-like charge order. The electron bubble phases of interest are identified by the conductivity minima located between these major features and adjacent satellite conductance peaks. For clarity, we adopt the conventional notation [11] $Na$ and $Nd$ (e.g., $4a,4d$) to label different bubble phases, where the integer $N$ indexes the specific spin-split Landau level branch, and the suffixes $a/d$ denote particle-hole conjugate states within that branch. To probe the evolution and thermal stability of these states, we track the change in magneto-conductivity at different temperatures. We find that the broad conductivity minima initially observed at base temperature gradually narrow and vanish at a certain point. This evolution marks the melting of electron-bubble phases, allowing the extraction of the solid-liquid transition temperature ($T_{m}$) by identifying peaks in the non-monotonic temperature dependence of the conductivity. We also observe that the thermal stability of bubble phases varies significantly with filling factor; $T_{m}$ is maximized at an optimal filling factor (indicated by the red triangle in Fig. 1b) and decreases away from this point, forming a characteristic dome-like phase diagram.
@@
As a large melting temperature indicates a more stable phase, we fixed the magnetic field at these filling factors and performed a detailed temperature-dependent conductivity measurement (Fig. 1c). The melting temperature ($T_{m}$) from conductivity peaks is extracted as a function of the partial filling factor ($\nu=\nu_{0}-2N$) at different Landau levels (Fig. 2, top panel). With increasing $\nu$ in a given Landau level, $T_{m}$ roughly follows a smooth decreasing trend, yet exhibits a discrete jump when transitioning to a higher Landau level. This behavior can be qualitatively explained by adapting the universal KTHNY transition picture to the bubble phase. If we treat the bubbles as classical objects with effective charge $Q=Me$ (where $M$ is the number of electrons per bubble), the KTHNY theory yields the melting criterion [18] $4\pi k_{B}T_{c}=0.283e^{2.5}M^{1.5}\sqrt{\frac{B\nu^{*}}{h}}$. Where $\nu^{*}=\nu\text{ mod}\,\frac{1}{2}$ is the partial filling factor defined as the decimal part of filling factor $\nu$. $k_{B}$ is the Boltzmann constant, $T_{c}$ is the critical temperature. This equation shows the transition temperature is a function of site electron occupation M, which is theoretically predicted to be dependent on Landau level $N$, with $M=N$ and $M=N+1$ being the most possible configurations. Assuming M=N under the classical picture, we calculate the expected temperature ratios between consecutive Landau levels and compare them with our experimental results. The predicted ratios of 1.49 (for $N=2\to 3$) and 1.07 (for $N=3\to 4$) show reasonable agreement with the observed values of 1.35 and 1.17, respectively.
@@
However, despite this simple classical picture capturing the qualitative inter-level jumps, it remains fundamentally inadequate that the absolute $T_{c}$ derived from it is severely overestimated. For example, for the N4a phase ($\nu=4.27$), the formula gives $T_{c}\sim 13K$, far above the measured SLT point $\sim 0.15K$. Moreover, the functional dependence within a Landau level deviates significantly from the prediction. Our data exhibits a nearly linear dependence of $T_{m}$ on the magnetic field, while classical theory predicts a square-root scaling ($\propto\sqrt{B}$).
@@
In conclusion, we perform bulk Corbino transport measurements that track the evolution of electron bubble phases across a wide range of Landau levels, and develop a theoretical framework that quantitatively captures their melting temperatures. Beyond identifying the conductivity and thus $T_{m}$ of a 2D electronic crystal, the comparison between experiment and theory allows us to extract the dislocation-core parameter $\alpha$ and the screening strength $\gamma$ in the relevant Landau level. Here $\alpha$ quantifies the relative importance of dislocation core energetics versus long-wavelength elasticity (i.e., the fragility of the electronic solid to fluctuation-driven phase transition), while $\gamma$ calibrates the effective interaction, thereby constraining the competing electronic phases as well as the energy scales of their collective modes and excitations. The experimentally obtained SLT boundary, therefore, provides a bulk constraint on the microscopic cost of creating topological defects and the dressing of interaction.
Our strategy can be extended to include disorder, anisotropy, finite thickness, and Landau-level mixing. More broadly, it provides a route to study the melting and phase competition in other 2D electronic crystals, including solids proximate to fractional quantum Hall liquids, where the fundamental effective particles are emergent quasiparticles (anyons). Understanding the existence of solid phases there is essential for studying the boundaries of topological order. It also applies to the recently observed generalized Wigner and multi-electron crystals in TMD moiré systems [30, 39, 21, 31], where the required energetics and elastic response can be computed from band-projected form factors and interactions, enabling a unified description of flat-band screening, interaction dressing, and the stability of charge-ordered phases in Chern-band systems.